# Exact Expressive Power of Transformers with Padding

### v1

@@ 5 Logd-Looped Padded Transformers Recognize Exactly $\mathsf{TC}^{d}$
Let $b(i)$ be the binary encoding of $i\in\mathbb{N}$ in some alphabet $\Sigma\supseteq\{0,1\}$. Let $\mathsf{R}$ be a class of languages. We say a transduction $f:\Sigma^{*}\to\Sigma^{*}$ is an $\mathsf{R}$ reduction if $\lvert f(w)\rvert$ is polynomial in $\lvert w\rvert$ and the language $R_{f}=\{(w,b(i),\sigma)\mid f(w)_{i}=\sigma\}$ is in $\mathsf{R}$ [^1] Here $f(w)_{i}$ denotes the $i$-th token of $f(w)$ if $i\leq\lvert f(w)\rvert$, and a special symbol $\square\not\in\Sigma$ otherwise.
Definition 9 recovers the standard notions of $\mathsf{FO}$ and $\mathsf{L}$ reductions. A transformer can be said to compute a reduction $f$ if it recognizes the language of triples $(w,i,\sigma)$ defined above. Since $\lvert\Sigma\rvert$ is finite and transformers can equally easily output a ‘token’ in $\Sigma$ instead of just 0/1, it is in fact more natural to require the transformer to compute a functional form of this language, namely compute $r_{f}(w,i)$ defined as $f(w)_{i}$. Our constructions work under both views, though the latter is often more natural and efficient. Formally:
@@
Proof.
Consider any language $L^{\prime}\in\mathsf{C}$. By the assumed completeness of $L$, there exists an $\mathsf{R}$ reduction $f$ that maps inputs of $L^{\prime}$ into inputs of $L$ such that $w\in L^{\prime}$ if and only if $f(w)\in L$. From the last precondition of the theorem, there exists a causally masked logd-depth padded transformer $T_{f}$ that recognizes the corresponding reduction language $R_{f}$ or, equivalently, computes the reduction function $r_{f}$ (cf. Definition 10). We will assume the latter, i.e., that $T_{f}$ computes $r_{f}$, though the construction can also be made to work if $T_{f}$ checks membership in $R_{f}$. Additionally, we also have from a precondition that there exists a causally masked logd-depth padded transformer $T_{L}$ that recognizes $L$.
Compute Reduction. Suppose $T_{f}\in\mathsf{AHAT}^{d}_{k}$ and let $n=\lvert w\rvert$. Then, on input $(w,i)$, where $i$ is represented in binary, $T_{f}$ uses $O(n^{k})$ padding tokens to compute $f(w)_{i}$, which we upper bound by $n^{k+1}$ for sufficiently large $n$ (for small $n$, we assume $T_{f}$ instead uses a finite lookup table (and no padding). There is a uniform $\mathsf{TC}^{0}$ circuit family that computes $\mathsf{bit}(i,j)$ from input $\phi(i),\phi(j)$, so, by Theorem 1, there exists a transformer $T_{\mathsf{bit}}$ that computes $\mathsf{bit}$ with $n^{b}$ padding tokens, for some $b$ and sufficiently large $n$ (again, for smaller $n$, we assume $T_{\mathsf{bit}}$ uses a finite lookup table). Since $\lvert f(w)\rvert$ is bounded by some polynomial $n^{c}$, we know that any $i\leq\lvert f(w)\rvert$ takes at most $c\log n$ bits to specify, which we upper bound by $c^{\prime}=n$ (since it’s unclear how a transformer would compute $\log n$).
We will construct $n^{c}$ blocks of padding tokens, each of size $B=c^{\prime}n^{b}+n^{k+1}$. Block $i$ will first compute the binary expansion of $i$ in its first $c^{\prime}n^{b}$ padding tokens using $T_{\mathsf{bit}}$, and then use the remaining $n^{k+1}$ padding tokens to consume $(w,i)$ using $T_{f}$ and return $r_{f}(w,i)$. At each token $t$ in block $i$, we compute $\phi(i)$ as $\phi(\lfloor t/B\rfloor)$, using the fact that we can compute integer division with a fixed block of transformer layers (Merrill and Sabharwal, 2024) and having computed $B$ as a function of $n$ in earlier layers (since it is in $\mathsf{TC}^{0}$). For $t^{\prime}\leq c^{\prime}$, token $t=t^{\prime}n^{b}$ computes $\phi(t^{\prime})$ via division similarly to $\phi(i)$ and then $\mathsf{bit}(i,t^{\prime})$. This recovers the binary representation of $i$ stored across tokens $t=t^{\prime}n^{b}$ in block $i$, for $1\leq t^{\prime}\leq c^{\prime}$. Finally, we apply a slightly modified $T_{f}$ over the block. Specifically, we add a new term to each attention head so that it only attends over the input tokens and some tokens within the current block: those satisfying $t=t^{\prime}n^{b}$ for some $1\leq t^{\prime}\leq c^{\prime}$ or $t>c^{\prime}n^{b}$: since $c^{\prime}=n$, these predicates are simple to check. As a result, block $i$ simulates $T_{f}$ over $w$ with $n^{k}$ padding. Thus, the final token of block $i$ computes $r_{f}(w,i)=f(w)_{i}$.
Combined with results in prior work (see proof below), it follows from Lemma 3 that padded log-depth transformers can recognize any language in $\mathsf{NL}$:
Theorem 2.
@@
Let $L$ be the graph connectivity problem, class $\mathsf{C}$ be $\mathsf{NL}$, and class $\mathsf{R}$ be $\mathsf{FO}$. We will show that the preconditions of Lemma 3 are met, from which it will follow that $\mathsf{NL}\subseteq\mathsf{AHAT}^{1}_{*}$. First, graph connectivity is known to be $\mathsf{NL}$-complete under $\mathsf{FO}$ reductions (Immerman, 1998). Second, Merrill and Sabharwal (2024) recently showed that mixed-masked log-depth transformers with cubic padding can recognize the graph connectivity problem $L$, i.e., $L\in\mathsf{mAHAT}^{1}_{*}$. From Proposition 1, it follows that $L\in\mathsf{AHAT}^{1}_{*}$. Finally, it follows from Theorem 1 that fixed-depth causally masked transformers can recognize languages in $\mathsf{FO}[\mathsf{M},\mathsf{bit}]$, and hence also languages in $\mathsf{FO}$. Such transformers can therefore compute $\mathsf{FO}$ reductions $f$ in the sense of Definition 10, i.e., compute the function $r_{f}(w,i)$ defined as $f(w)_{i}$, the $i$-th bit of $f(w)$. Thus, Lemma 3 applies. ∎
Since $\mathsf{L}$ reductions are in $\mathsf{NL}$, we can bootstrap this result to obtain the following stronger result. For this, we will leverage completeness of the problem of evaluating a given logd depth circuit, formalized in Definition 11 (Appendix B) as the $\mathsf{TC}^{d}$ circuit evaluation problem. We first note the following extension of a known result about fixed-depth transformers (see Appendix A for a proof):
@@
For any $d\geq 1$, $\mathsf{AHAT}^{d}_{*}\subseteq\mathsf{L}\textrm{-uniform}\ \mathsf{TC}^{d}$.
In contrast to Theorem 1, Lemma 4 applies with $\mathsf{L}$ uniformity instead of $\mathsf{FO}$ uniformity. We will show the other direction of Lemma 4 to establish $\mathsf{L}$-uniform $\mathsf{TC}^{d}$ as an exact characterization of $\mathsf{AHAT}^{d}_{*}$:
@@
For any $d\geq 1$, $\mathsf{AHAT}^{d}_{*}=\mathsf{L}$-uniform $\mathsf{TC}^{d}$. [^2] Theorem 3 generalizes to $\mathsf{NL}$-uniform $\mathsf{TC}^{d}=\mathsf{AHAT}^{d}_{*}$, which implies $\mathsf{NL}$-uniform $\mathsf{TC}^{d}=\mathsf{L}$-uniform $\mathsf{TC}^{d}$. At first glance, it might appear suspicious that we can prove this complexity result through our analysis of transformers. However, we verify in Appendix C via a simpler proof that this equivalence indeed holds.

### v2

@@ 5 Logd-Looped Padded Transformers Recognize Exactly $\mathsf{FO}\textrm{-uniform}\ \mathsf{TC}^{d}$
Let $b(i)$ be the binary encoding of $i\in\mathbb{N}$ in some alphabet $\Sigma\supseteq\{0,1\}$. Let $\mathsf{R}$ be a class of languages. We say a transduction $f:\Sigma^{*}\to\Sigma^{*}$ is an $\mathsf{R}$ reduction if $\lvert f(w)\rvert$ is polynomial in $\lvert w\rvert$ and the language $R_{f}=\{(w,b(i),\sigma)\mid f(w)_{i}=\sigma\}$ is in $\mathsf{R}$. [^2] Here $f(w)_{i}$ denotes the $i$-th token of $f(w)$ if $i\leq\lvert f(w)\rvert$, and a special symbol $\square\not\in\Sigma$ otherwise.
Definition˜9 recovers the standard notions of $\mathsf{FO}$ and $\mathsf{L}$ reductions. A transformer can be said to compute a reduction $f$ if it recognizes the language of triples $(w,i,\sigma)$ defined above. Since $\lvert\Sigma\rvert$ is finite and transformers can equally easily output a ‘token’ in $\Sigma$ instead of just 0/1, it is in fact more natural to require the transformer to compute a functional form of this language, namely compute $r_{f}(w,i)$ defined as $f(w)_{i}$. Our constructions work under both views, though the latter is often more natural and efficient. Formally:
@@
Proof Sketch.
Consider any language $L^{\prime}\in\mathsf{C}$. By the assumed completeness of $L$, there exists an $\mathsf{R}$ reduction $f$ that maps inputs of $L^{\prime}$ into inputs of $L$ such that $w\in L^{\prime}$ if and only if $f(w)\in L$. From the last precondition of the theorem, there exists a causally masked logd-depth padded transformer $T_{f}$ that recognizes the corresponding reduction language $R_{f}$ or, equivalently, computes the reduction function $r_{f}$ (cf. Definition˜10). We will assume the latter, i.e., that $T_{f}$ computes $r_{f}$, though the construction can also be made to work if $T_{f}$ checks membership in $R_{f}$. Additionally, we also have from a precondition that there exists a causally masked logd-depth padded transformer $T_{L}$ that recognizes $L$.
The idea is to “stack” $T_{L}$ on top of $T_{f}$ to obtain a logd-depth padded transformer $T$ that recognizes $L^{\prime}$. Intuitively, given an input $w$, the first set of layers of $T$ will compute $f(w)$ tokenwise, by computing $r(w,i)=f(w)_{i}$ for every $i$ in parallel. To this end, we will essentially make $\lvert f(w)\rvert$ copies of the padding tokens needed by $T_{f}$ and perform the computation of $T_{f}$ independently for each $f(w)_{i}$. The second and final set of layers of $T$ will then check whether the string $f(w)$ produced by the first set of layers is in $L$, which will hold if and only if $w\in L^{\prime}$. See full proof in Appendix˜C. ∎
Combined with results in prior work, it follows from Lemma˜3 (see proof below) that padded log-depth transformers can recognize any language in $\mathsf{NL}$:
Lemma 4.
@@
Let $L$ be the graph connectivity problem, class $\mathsf{C}$ be $\mathsf{NL}$, and class $\mathsf{R}$ be $\mathsf{FO}$. We will show that the preconditions of Lemma˜3 are met, from which it will follow that $\mathsf{NL}\subseteq\mathsf{AHAT}^{1}_{*}$. First, graph connectivity is known to be $\mathsf{NL}$-complete under $\mathsf{FO}$ reductions (immerman-1998-descriptive). Second, merrill-2025-little recently showed that mixed-masked log-depth transformers with cubic padding can recognize the graph connectivity problem $L$, i.e., $L\in\mathsf{mAHAT}^{1}_{*}$. From Proposition˜1, it follows that $L\in\mathsf{AHAT}^{1}_{*}$. Finally, it follows from Theorem˜1 that fixed-depth causally masked transformers can recognize languages in $\mathsf{FO}[\mathsf{M},\mathsf{bit}]$, and hence also languages in $\mathsf{FO}$. Such transformers can therefore compute $\mathsf{FO}$ reductions $f$ in the sense of Definition˜10, i.e., compute the function $r_{f}(w,i)$ defined as $f(w)_{i}$, the $i$-th bit of $f(w)$. Thus, Lemma˜3 applies. ∎
Since $\mathsf{L}$ reductions are in $\mathsf{NL}$, we can bootstrap this result to obtain the following stronger result. For this, we will leverage the notion of reductions and the completeness of the problem of evaluating a given “wide” logd depth circuit, formalized in Definition˜15 (Appendix˜D).
@@
For $d\geq 0$, $\mathsf{FO}\textrm{-uniform}\ \mathsf{TC}^{d}\subseteq\mathsf{AHAT}^{d}_{*}$.
Proof.
We will apply Lemma˜3 with the wide-$\mathsf{TC}^{d}$ circuit evaluation problem (Appendix˜D) as $L$, $\mathsf{FO}\textrm{-uniform}\ \mathsf{TC}^{d}$ as class $\mathsf{C}$, and $\mathsf{L}$ as class $\mathsf{R}$. We next argue that the preconditions of Lemma˜3 are met, which will finish the proof. First, ˜13.1 (Section˜D.2) shows that logd-depth looped transformers (without padding) can solve the wide-$\mathsf{TC}^{d}$ circuit evaluation problem, i.e., $L\in\mathsf{AHAT}^{d}_{0}$. Second, ˜14.2 (Section˜D.3) shows that $L$ is complete for $\mathsf{FO}\textrm{-uniform}\ \mathsf{TC}^{d}$ under $\mathsf{L}$ reductions. Finally, Lemma˜4 implies that logd-depth transformers for $d\geq 1$ can recognize any language in $\mathsf{L}$, and thus compute any $\mathsf{L}$ reduction in the sense of Definition˜10. Applying Lemma˜3, we conclude that $\mathsf{FO}\textrm{-uniform}\ \mathsf{TC}^{d}\subseteq\mathsf{AHAT}^{d}_{*}$. ∎
The proof heavily leverages the fact that wide-$\mathsf{TC}^{d}$ circuit evaluation $\mathsf{TC}^{d}$-complete, which we show in Section˜D.1. To our knowledge, formalizing this $\mathsf{TC}^{d}$-complete problem (or, in fact, any natural $\mathsf{TC}^{d}$-complete problem) is a novel contribution. We next note the following extension of a known result about fixed-depth transformers. See Appendix˜C for a proof, which leverages the fact that recurrent composition of poly-size $\mathsf{FO}\textrm{-uniform}$ circuit families remains $\mathsf{FO}\textrm{-uniform}$ (Lemma˜9, Section˜B.1):
Lemma 6.
For $d\geq 1$, $\mathsf{AHAT}^{d}_{*}\subseteq\mathsf{FO}\textrm{-uniform}\ \mathsf{TC}^{d}$.
Combining Lemmas˜6 and 5, we obtain an exact characterization of $\mathsf{AHAT}^{d}_{*}$ for $d\geq 1$:
@@
For any $d\geq 1$, $\mathsf{AHAT}^{d}_{*}=\mathsf{FO}\textrm{-uniform}\ \mathsf{TC}^{d}$.
Taking the union over all $d$, we obtain $\bigcup_{d=0}^{\infty}\mathsf{AHAT}^{d}_{*}$ as the class of languages recognized by polylogarithmic-looped padded transformers. Theorems˜1 and 2 imply that this class is the same as $\bigcup_{d=0}^{\infty}\mathsf{FO}\textrm{-uniform}\ \mathsf{TC}^{d}$, which in turn is the same as the class $\mathsf{FO}\textrm{-uniform}\ \mathsf{NC}$, where $\mathsf{NC}$ is the class of all “parallelizable” languages—those recognized by polylogarithmic depth (and polynomial size) circuit families. We therefore have:
Corollary 2.1.
Polylog-looped poly-padded transformers recognize exactly $\mathsf{FO}\textrm{-uniform}\ \mathsf{NC}$.
As before, we note that the arguments leading to Theorems˜2 and 2.1 work for both logarithmic and polynomial precision (cf. Appendix˜A) transformers. This shows that these two levels of precision lead to identical power for polylog-looped polynomially-padded transformers.
6 Uniformity Collapse for Circuit Classes
Our results on looped transformers involved substantial analysis of uniform polylogarithmic-depth circuit classes. This analysis led us to prove a novel result about uniform circuit families, which both slightly strengthens our results about looped transfomers and may be of independent interest.
The result concerns the strength of different uniformity conditions for circuit families, characterizing conditions under which variants of circuit classes with different uniformity conditions collapse to express the same class of languages (Proposition˜6 in Appendix˜E). Applying this general result to (wide) $\mathsf{TC}^{d}$ circuits, we show that for $d\geq 1$, both $\mathsf{NL}\textrm{-uniformity}$ and $\mathsf{L}\textrm{-uniformity}$ collapse to the weaker notion of $\mathsf{FO}\textrm{-uniformity}$ for both $\mathsf{AC}^{d}$ and $\mathsf{TC}^{d}$ circuits:
Theorem 3 (Uniformity Collapse).
For any $d\geq 1$, the following equivalences hold:
$\mathsf{FO}\textrm{-uniform}\ \mathsf{AC}^{d}$ $=\mathsf{L}\textrm{-uniform}\ \mathsf{AC}^{d}=\mathsf{NL}\textrm{-uniform}\ \mathsf{AC}^{d}$
$\mathsf{FO}\textrm{-uniform}\ \mathsf{TC}^{d}$ $=\mathsf{L}\textrm{-uniform}\ \mathsf{TC}^{d}=\mathsf{NL}\textrm{-uniform}\ \mathsf{TC}^{d}.$
The key idea, formalized in Appendix˜E, is to leverage the fact that $\mathsf{NL}$ and $\mathsf{L}$ themselves are in $\mathsf{FO}\textrm{-uniform}\ \mathsf{AC}^{d}$ for $d\geq 1$, and therefore the $\mathsf{NL}$ or $\mathsf{L}$ machine that builds a circuit family can itself be simulated in $\mathsf{FO}\textrm{-uniform}\ \mathsf{AC}^{d}$ as well. Thus all one needs to do is compose this “circuit building” circuit family with a “circuit evaluation” circuit family. To this end, we show that functions computed by $\mathsf{FO}\textrm{-uniform}$ circuit families are closed under fixed compositions (Proposition˜3).
Theorem˜3 implies that, for $d\geq 1$, $\mathsf{AHAT}^{d}_{*}$ recognizes not just $\mathsf{FO}\textrm{-uniform}\ \mathsf{TC}^{d}$ but also $\mathsf{L}\textrm{-uniform}\ \mathsf{TC}^{d}$ because these classes are the same.
7 Conclusion
Our results in this work give a precise theoretical understanding of how padding and looping—two ways to dynamically expand the computational resources of a transformer at inference time—increase the expressive power of transformers. Padding expands the circuit width of transformers, allowing them to resolve logical formulas over more variables. As a consequence of this, polynomially padded transformers can recognize exactly $\mathsf{TC}^{0}$, which was previously known only as an upper bound (merrill-sabharwal-2023-parallelism; pfau-2024-think). In contrast, looping increases the depth of transformers. Applying looping on top of padding, we extended our result to show that logd-depth, padded transformers recognize exactly $\mathsf{TC}^{d}$. This means that transformers with polynomial padding and polylogarithmic looping converge to recognizing $\mathsf{NC}$, the largest class of problems that can be solved with parallel computation. In contrast, transformers with CoT have greater expressive power under standard complexity conjectures (merrill2023cot), but suffer from slow sequential decoding. Thus, while looping and padding are not as powerful as CoT, our results suggest they are quite effective ways to expand transformers’ expressive power while preserving parallelism.
Several interesting open questions remain from this work. On the theoretical side, it would be valuable to develop a more finegrained characterization of $\mathsf{AHAT}^{d}_{k}$, i.e., looped transformers where the padding is at most $\mathrm{O}(n^{k})$ for some fixed $k$. On the empirical side, while looped and padded transformers have both been explored already to an extent, it would be interesting to see whether these approaches could be successfully integrated to improve the performance of transformers on hard reasoning tasks. Further developing these approaches could ultimately lead to inference-time compute methods based on padding and looping that increase transformers’ expressive power without sacrificing parallelism, providing a more efficient alternative to CoT for solving moderately parallelizable problems.
Limitations
Practical Concerns for Looped Transformers with Padding. In this paper, we showed looped transformers with padding are quite expressive. However, the degree to which transformers can learn to use looped layers is an important open question. In particular, practical details might be important here, such as an appropriate parameterization where features can be learned in later layers (dey2025completep). Another practical caveat is that, while padding is efficiently parallelizable, it will extend the context length, incurring more memory overhead for the forward pass.
Transformer Model. Here we have assumed AHATs here rather than softmax-attention transformers (SMATs). For any fixed maximum context length, it is possible scale the temperature of SMAT heads to arbitrarily approximate AHAT heads, but for unbounded context lengths, SMATs may not be able to simulate AHATs. We thus view the AHAT as mild simplification of the SMAT that abstracts away issues with soft attention for simulating hard attention over very long contexts. It would also be interesting to better understand the necessity of the masked pre-norm assumption.
Acknowledgments
We appreciate discussions with Selim Jerad, Andy Yang, Michael Cadhilac, and attendees of the Formal Languages and Neural Networks (FLaNN) seminar. This project was supported by the National Science Foundation (NSF) through award 1922658 and WM’s NSF Graduate Research Fellowship. WM was also supported by a Two Sigma PhD Fellowship and the Allen Institute for AI.
Appendix A Datatype Assumptions
We adapt the $p$-precise datatype model from merrill-2025-little. We encode scalars as strings in $\{0,1\}^{p}$, where $p$ can be a function of $n$. If $p$ depends on $n$, activations, but not models parameters, can depend on $n$. A datatype $\mathbb{D}_{p}$ assigns a numerical semantics for each string in $\{0,1\}^{p}$. For $x\in\mathbb{R}$, let $x]_{\mathbb{D}_{p}}$ be $x$ rounded into $\mathbb{D}_{p}$, i.e., the bitstring whose numerical value in $\mathbb{D}_{p}$ is closest to $x$ (breaking ties in favor of the higher value). We define our datatype $\mathbb{D}_{p}$ to satisfy the following:
Definition 11 ($p$-Precise Operations).
Let $f:\mathbb{R}^{k}\to\mathbb{R}$ be an operation with $p$-precision realization $\tilde{f}:\mathbb{D}_{p}^{k}\to\mathbb{D}_{p}$. We say $\tilde{f}$ is $p$-precise if, for any $x_{1},\ldots,x_{k}\in\mathbb{R}$ exactly representable in $\mathbb{D}_{p}$,
$f(x_{1},\ldots,x_{k})]_{\mathbb{D}_{p}}=\tilde{f}([x_{1}]_{\mathbb{D}_{p}},\ldots,[x_{k}]_{\mathbb{D}_{p}}).$
To apply Definition˜11, we view the summation in attention heads as an $n$-ary operation. We also view layer-norm as a single operation from $\mathbb{R}^{m}\to\mathbb{R}^{m}$. Lastly, we assume that these operations in the computation graph are defined in the standard way and are thus computable in $\mathsf{TC}^{0}$ [merrill-sabharwal-2023-parallelism, merrill2023logic].
We consider two natural instantiations of $\mathbb{D}_{p}$ in the main text: log precision, with $p=c\log n$ [merrill-sabharwal-2023-parallelism], and polynomial precision, with $p=n^{c}$ [chiang2025transformers], for some fixed $c>0$. All our results go through with either datatype, showing their equivalence for polylogarithmically-looped polynomially-padded transformers.
Appendix B Uniformity of Circuit Classes
To make our arguments rigorous for highly uniform (i.e., $\mathsf{FO}\textrm{-uniform}$) circuit families, it will be necessary to work with a detailed definition of uniformity. For this, we start with a standard formal definition of the connection language describing the gates and wires of a circuit family. Here we allow each circuit to have more than one (numbered) output gates.
Definition 12 (Connection Language; cf. vollmer1999introduction).
Define the connection language for a circuit family $\mathcal{C}=\{C_{n}\}_{n=0}^{\infty}$ as $L_{\mathcal{C}}=L_{\textrm{gate}}\cup L_{\textrm{wire}}$ where
[…]
$L_{\textrm{wire}}$ $=\{a^{n}ij\mid\textrm{$C_{n}$ has a wire $i\to j$}\}$
and $i,j\in\mathbb{Z}_{\geq 0}$, $g$ is a gate type, and $k,\ell\in\mathbb{Z}_{\geq-1}$. Here $i,j,k,\ell$ are represented using exactly $c\log n$ bits for some fixed $c\in\mathbb{Z}_{\geq 1}$.
We will assume that gates in all circuits are numbered so that there is a block of input gates, followed by a block of intermediate gates, followed by a block of output gates. Thus, there exist specific gate thresholds that denote the boundaries for input and output gates.
Definition 13 (Generalized Uniformity).
A family of circuits $\mathcal{C}=\{C_{n}\}_{n=0}^{\infty}$ is $\mathsf{A}\textrm{-uniform}$ if $L_{\mathcal{C}}\in\mathsf{A}$.
When $\mathsf{A}=\mathsf{L}$, this generalized uniformity notion is equivalent to the standard definition of uniformity in terms of being able to serialize $C_{n}$ as a function of $1^{n}$:
Proposition 2.
$\mathcal{C}=\{C_{n}\}_{n=0}^{\infty}$ is $\mathsf{L}\textrm{-uniform}$ if and only if $1^{n}\mapsto\langle C_{n}\rangle$ is computable in log space.